Operations Report – 26 April 2018

SOL: 4

Name of person filing report: David Attig

Non-nominal systems: Com radio #6, Sim suits #10, 2, 5, 4, 6.

Notes on non-nominal systems: Status unchanged since yesterday.

Generator (hours run): 9:00 pm to 9:00 am.

Solar— SOC 80 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 550 gallons.

Water (static) – 305 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134126.7

Toilet tank emptied: yes
Date last emptied: April 24

ATV’s (Honda, 300, 350.1, 350.2, 350.3): No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 is out of service, due to a previous issue.

Deimos rover used: Yes
Hours: 124.8
Beginning charge: 100%
Ending charge: 94 %
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes
Hours: 35.0
Beginning charge: 100 %
Ending charge: 91 %
Currently charging: Yes

Opportunity rover used: Out Of Service
Hours: 31
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: 91 %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: I took pans out of broiler underneath
oven; It is an older style gas stove, with a broiler (not a drawer)
underneath the oven. Too many pans there could block airflow for oven
burner.

Summary of internet: As of 9:30 AM, we have 40% of our internet
allowance remaining. The funny thing is that our internet while we
actually do have some percentage remaining doesn’t seem any faster
than at night when we are hard throttled.

Summary of suits and radios: All one piece suits are working (and
were used today). Radios are nominal.

Summary of Hab operations: Supported two (4 hour) EVAs.

Summary of GreenHab operations: Picked some plants for lunch today.

Summary of ScienceDome operations: Samples placed in incubator.

Summary of RAM operations: Morning EVA team brought materials in from
RAM upon returning. New radios are in the RAM airlock for us.

Summary of health and safety issues: None to report.

Operations Report April 25th

Crew 193 Operations Report 25Apr2018

SOL: 3

Name of person filing report: David Attig

Non-nominal systems: Com radio #6, Sim suits #10, 2, 5, 4, 6.

Notes on non-nominal systems: Radio #06 (newer style) only lasted 1 hour on EVA with a freshly charged battery. Radio will not be used for future EVAs. Sim suit 10 still has the badly scratched visor, suits 2, 5, 6 need maintenance on straps attaching neck ring to pack, and suit 4 has a broken charge connector.

Generator (time run): 2000-0715

Solar— SOC 95 % (Before generator is run at night)

Diesel – 75 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 3500 gallons.

Water (storage) – 550 gallons

Water (static) – 350 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134078.1

Toilet tank emptied: no
Date last emptied: April 24

ATV’s (Honda, 300, 350.1, 350.2, 350.3):
Honda Used?: Yes
300 Used? No
350.1 Used? Yes
350.2 Used? No
350.3 Used? No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 1/4 Gallons

# Hours the ATVs were Used today: 1.2

Notes on ATVs: None

Deimos rover used: No
Hours: 124.0
Beginning charge: 100%
Ending charge: 100 %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 33.7
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Opportunity rover used: Out Of Service
Hours: 31
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: 23
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: We would like the sim suit supply box dropped off in our airlock, so we can work on the two-piece suits.

Summary of internet: Appears to be working.

Summary of suits and radios: Radio 06 appears to transmit a beep continuously, draining the battery. Suits have some issues; see email titled “Details about EVA Suits”.

Summary of Hab operations: supported two EVAs and investigated suit status. Sol Summary report for more details.

Summary of GreenHab operations: plants are doing well, see GreenHab report for more details.

Summary of ScienceDome operations: Progress was made on dishwasher experiment, see Science Report for more details.

Summary of RAM operations: Placed tools in RAM for future use, took boards apart and painted some to use for an MDRS sign.

Summary of health and safety issues: Need to work on better balance between working in RAM and not breaking sim; see email “A few notes – about EVA duration” for details.

Questions, concerns and requests to Mission Support: A few requests, listed in Sol Summary.

Operations Report – April 24th

Crew 193 Operations Report 24Apr2018

SOL: 2

Name of person filing report: David Attig

Non-nominal systems: One piece EVA suit #10 and two piece EVA suit # 6 have issues

Notes on non-nominal systems: One two piece suit (6) has straps broken holding on neck ring, Visor on suit 10 is heavily scratched and deemed unuseable. See Sol Summary for more details.

Generator (hours run): 12

Solar— SOC 86 % (Before generator is run at night)

Diesel – 75 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0.75 gallons.

Water (trailer) – 0 gallons.

Water (storage) – 550 gallons

Water (static) – 410 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134008.9

Toilet tank emptied: yes
Date last emptied: April 24

ATV’s (Honda, 300, 350.1, 350.2, 350.3):
Honda Used?: Yes
300 Used? Yes
350.1 Used? No
350.2 Used? Yes
350.3 Used? No

ATV Oil Checked?: Yes

ATV Oil Added?: No

ATV Fuel Used: less than 1 Gallon

# Hours the ATVs were Used today: 0.5

Notes on ATVs: With 350.2, throttle cable gets pulled when handlebar is moved all the way to the left. Suggest not using this ATV for future (provided sufficient fuel in others).

Deimos rover used: Yes
Hours: 124.0
Beginning charge: 99%
Ending charge: 98 %
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 33.7
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Opportunity rover used: Out Of Service
Hours: 31
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: 23
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments:

Summary of internet: By 8AM again today, we have used half of our allotted daily internet. Is there a way to monitor our internet over time without constantly logging on to check it? That way we could monitor usage and know if there is a peak time that it is being used.

Summary of suits and radios: Older style (silver front) radios appear to not transmit well enough to use in EVA. One two piece suit broke during EVA today, and 1 one piece suit was deemed unuseable due to scratches in the visor.

Summary of Hab operations: See Sol Summary report.

Summary of GreenHab operations: See GreenHab report

Summary of ScienceDome operations: See Science report.

Summary of RAM operations: We cut a board to use for a new sign in front of MDRS.

Summary of health and safety issues: None to report.

Questions, concerns and requests to Mission Support: See Sol Summary.

Operations Report – April 23rd

Crew 193 Operations Report 23Apr2018

SOL: 1

Name of person filing report: David Attig

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): 14

Solar— SOC 92 % (Before generator is run at night)

Diesel – 75 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 1 gallons.

Water (trailer) – 0 gallons.

Water (storage) – 550 gallons

Water (static) – 470 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 133949.8

Toilet tank emptied: no
Date last emptied: April 21

ATV’s (Honda, 300, 350.1, 350.2, 350.3):
Honda Used?: Yes
300 Used? No
350.1 Used? Yes
350.2 Used? Yes
350.3 Used? Yes

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0.3

Notes on ATVs: None

Deimos rover used: No
Hours: 123.5
Beginning charge: 100%
Ending charge: 100 %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 33.7
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Opportunity rover used: Out Of Service
Hours: 31
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: 23
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: This morning we had an issue with the readings on the solar SoC before we turned off the generator. Systems seem to be nominal now.

Summary of internet: Nominal. This morning the internet was very slow, while we still had 49% of our

Summary of suits and radios: All one piece suits are working (and were used today). Radios are nominal.

Summary of Hab operations: See Sol Summary report.

Summary of GreenHab operations: See GreenHab report

Summary of ScienceDome operations: 001 constructed dishwasher experiment. While checking SoC, I noticed one of the nuts had fallen from lock. It has been replaced, but is not fully tightened as that would interfere with Lock operations. We will keep an eye on it, suggest a longer bolt with two nuts jammed together in future upgrades.

Summary of RAM operations: Sign marking area of our settlement was in RAM. Old sign was repaired and moved outside. New sign material is in RAM awaiting construction. Door was left open, engineering EVA was performed to close it.

Summary of health and safety issues: None to report.

Questions, concerns and requests to Mission Support: Nothing yet.

Operations Report – April 21st

Crew 192 Operations Report 20Apr2018

SOL 13

Name of person filing report: Richard Blakeman

Non-nominal systems: Upper deck faucet leak continues.

Generator (hours run): 12 hours

Solar ‒ SOC 100%

Diesel – 75% fuel

Propane – 60% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 25% gas left in one can.

Water (trailer) – Empty water tank secure on Trailer.
Water (static) – 500 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133860.0

Toilet tank emptied: Yes. Last emptied today, April 21.

ATV’s Used: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: 0

Notes on ATVs: All working and ready to use.

Deimos rover used: No (decommissioned)
Hours: 122.3
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 32.8 (corrected)
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Opportunity rover used: No
Hours: 30.4
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 27.4
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? Yes, dinner with incoming crew and water runs.

General notes and comments: Last minute cleaning and tasks to prepare hab for incoming crew.

Summary of internet: 0MB at 1900.

Summary of suits and radios: All suits and radios cleaned, charged, and prepped for incoming crew.

Summary of Hab operations: Worked through departure checklist.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: None

Summary of RAM operations: None

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: None

Operations Report – April 20th

Crew 192 Operations Report 20Apr2018

SOL 12

Name of person filing report: Richard Blakeman

Non-nominal systems: Upper deck faucet, hot water now has to be turned off when not in use because it leaks too much, causing water loss and constant noise from the pump.

Generator (hours run): 12 hours

Solar ‒ SOC 87%

Diesel – 75% fuel

Propane – 60% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 25% gas left in one can.

Water (trailer) – Empty water tank secure on Trailer.
Water (static) – 550 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133825.0

Toilet tank emptied: Yes. Last emptied today, April 20.

ATV’s Used: Blue ATVs 1, 2, 3, and 4

Oil Added? No but checked and OK last time road.

ATV Fuel Used: 1, 2, and 3 at 75%, 4 at 80%

# Hours the ATVs were Used today: 6

Notes on ATVs: ATVs fine, oil checked.

Deimos rover used: No
Hours: 122.3
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 32.8 (corrected)
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Opportunity rover used: No
Hours: 30.4
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 27.4
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Cleaned the hab from top to bottom to prepare for incoming crew.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, two one-piece suits used, radios working fine. Comms headpieces all cleaned and prepared for incoming crew.

Summary of Hab operations: Cleaning and organizing for Crew 193.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: Last cleaning check.

Summary of RAM operations: None

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: None

Operations Report – April 19th

Crew 192 Operations Report 19Apr2018

SOL 11

Name of person filing report: Richard Blakeman

Non-nominal systems: Hot water faucet in upper deck of hab, cleaned out water pump for water tank.

Generator (hours run): 12 hours

Solar ‒ SOC 87%

Diesel – 75% fuel

Propane – 61% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – All ATVs filled 100%, 25% left in one can.

Water (trailer) – Empty water tank secure ob Trailer.
Water (static) – 690 gallons (two external tanks)
Trailer to Static Pump used – Yes
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133755.8

Toilet tank emptied: No. Last emptied April 16.

ATV’s Used: None

Oil Added? No but checked and OK last time road.

ATV Fuel Used: 0% today

# Hours the ATVs were Used today: 0

Notes on ATVs: ATVs fine, oil checked, gas full.
Assumption on percentage …untenable to check
Deimos rover used: No
Hours: 121.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 32.8 (corrected)
Beginning charge: 100%
Ending charge: N/A
Currently charging: Yes

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: N/A
Currently charging: Yes

Curiosity rover used: No
Hours: 27.4
Beginning charge: 100%
Ending charge: N/A
Currently charging: No

HabCar used and why, where? No

General notes and comments: A great EVA was followed by cleaning and preparing the hab for Crew 193.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, three one-piece suits used, radios working fine. Inventoried and cleaned 30 ear pieces not including 5 our crew has.

Summary of Hab operations: Cleaning and organizing for Crew 193.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: Thoroughly cleaned and prepared for next crew.

Summary of RAM operations: Trailer water tank moved out of RAM and back onto trailer and secured. Doctor requests water test kit before use. Cleaned, swept and mopped RAM…secured doors.

Summary of health and safety issues: Crew was very tired after working on water tank, so afternoon EVA was changed to an engineering EVA.

Questions, concerns, and requests to Mission Support: Fuel for ATVs

Operations Report – April 18th

Crew 192 Operations Report 18Apr2018

SOL 10

Name of person filing report: Victoria Varone

Non-nominal systems: None

Generator (hours run): 12 hours

Solar ‒ SOC 66%

Diesel – 75% fuel

Propane – 62% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – All ATVs filled 100%, 25% left in one can, requesting gas refill.

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 690 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133669.6

Toilet tank emptied: No. Last emptied April 16.

ATV’s Used: None

Oil Added? No but checked and OK.

ATV Fuel Used: 0% today

# Hours the ATVs were Used today: 0

Notes on ATVs: ATVs fine, oil checked, gas filled yesterday.

Deimos rover used: Yes (out of commission per Shannon)
Hours: 121.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 32.8 (corrected)
Beginning charge: 100%
Ending charge: 68%
Currently charging: Yes

Opportunity rover used: yes
Hours: 29.8
Beginning charge: 100%
Ending charge: 81%
Currently charging: Yes

Curiosity rover used: Yes
Hours: 27.4
Beginning charge: 100%
Ending charge: 67%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Some great EVAs on a beautiful day.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, three one-piece suits used, radios working fine.

Summary of Hab operations: Cooked, cleaned, and organized the hab.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: None

Summary of RAM operations: Trailer water tank remains in the RAM to air out after being cleaned for the final time yesterday, will be checked tomorrow.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: – Hab upper deck faucet hot water shut due to increasing leaks.

– Fridge got hot last night despite being closed and plugged in properly, was unplugged and plugged back and is working fine again.

– Steps in front of rear airlock were secured again after become a bit loose after the first fix.

Operations Report – April 17th

Crew 192 Operations Report 17Apr2018

SOL 9

Name of person filing report: Victoria Varone

Non-nominal systems: None

Generator (hours run): 12 hours

Solar ‒ SOC 87%

Diesel – 75% fuel

Propane – 62% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – All ATVs filled 100%, 25% left in one can, requesting gas refill.

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 730 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133616.7

Toilet tank emptied: No. Emptied yesterday, April 16.

ATV’s Used: Red ATV, Blue ATV’s #1, #2, and #3

Oil Added? No but checked and OK.

ATV Fuel Used: 20%

# Hours the ATVs were Used today: 6

Notes on ATVs: All ATVs working well, oil checked, gas filled today.

Deimos rover used: No
Hours: 120
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 31.2 (corrected)
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.9
Beginning charge: 91%
Ending charge: 91%
Currently charging: Yes

HabCar used and why, where? No

General notes and comments: Crew is doing well, trying to make the best use of time for EVAs in the midst of constant strong winds.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, three two-piece suits used, radios working fine. Suit #4 charging port is faulty.

Summary of Hab operations: Cleaned and organized hab.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: None

Summary of RAM operations: Trailer water was cleaned for final time today, remains in the RAM while strong winds persist.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: Doctor Narayanamoorthi recommends biological and chemical testing of water in trailer water tank before clearance for crew use.

Faucet in upper deck of hab was fixed as best we could but it continues to slowly leak.

Operations Report – April 16th

Crew 192 Operations Report 16Apr2018

SOL 8

Name of person filing report: Richard Blakeman

Non-nominal systems: None

Generator (hours run): 12 hours

Solar ‒ SOC 78%

Diesel – 70% fuel

Propane – 63% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Approximately 4 gallons left in one and the other is empty

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 780 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133570.4

Toilet tank emptied: Yes. Emptied today, April 16.

ATV’s Used: None

Oil Added? No but checked and OK.

ATV Fuel Used: 5%

# Hours the ATVs were Used today: 0.5

Notes on ATVs: All working well and ready for use, gas filled up yesterday.

Deimos rover used: Yes
Hours: 120
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 38.2
Beginning charge: 100%
Ending charge: 98%
Currently charging: Yes

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.9
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Crew is doing well, though the bad weather is hampering our EVAs and astronomical activities.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, 3 two-piece suits used, radios working fine.

Summary of Hab operations: Finished removing hab skirt as requested, kept a close eye on all hab systems as strong winds blew most of the day.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: None

Summary of RAM operations: Trailer water remains in the RAM for continued cleaning of embedded sediment.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: Requesting trash pickup.